OpenStudy (anonymous):

Can anyone tell me how to get the answer to 3x/24 (its a fraction) will medal and fan

jabez177 (jabez177):

You have to get x alone on one side.

OpenStudy (anonymous):

okay so you do x=3/24

OpenStudy (anonymous):

@dr0zier99

jabez177 (jabez177):

Okay, the answer is \[-\frac{ x }{ 8 }\]

OpenStudy (dr0zier99):

yes

OpenStudy (anonymous):

Oh ok i got it now thx anyway

OpenStudy (anonymous):

@jabez177 how did you get -

jabez177 (jabez177):

Because it's a negative question in the beginning and when you multiple a negative it stays negative or something like that. :P I forgot a little.

jabez177 (jabez177):

*Multiply

OpenStudy (anonymous):

but i didn't say it was negative its positive so is the answer just x/24?

jabez177 (jabez177):

Ohhhhh!! I'm so sorry!! I didn't see it! Let me calculate that again!

OpenStudy (anonymous):

Oh its ok no problem!

jabez177 (jabez177):

Okay, it's just \[\frac{ x }{ 8 }\] No negative.

OpenStudy (anonymous):

gotcha thanks

jabez177 (jabez177):

:)